# Integrating Canonical Neural Units and Multi-Scale Training for Handwritten Text Recognition

### v1

@@ Abstract
The segmentation-free research efforts for addressing handwritten text recognition can be divided into three categories: connectionist temporal classification (CTC), hidden Markov model and encoder–decoder methods. In this paper, inspired by the above three modeling methods, we propose a new recognition network by using a novel three-dimensional (3D) attention module and global-local context information. Based on the feature maps of the last convolutional layer, a series of 3D blocks with different resolutions are split. Then, these 3D blocks are fed into the 3D attention module to generate sequential visual features. Finally, by integrating the visual features and the corresponding global-local context features, a well-designed representation can be obtained. Main canonical neural units including attention mechanisms, fully-connected layer, recurrent unit and convolutional layer are efficiently organized into a network and can be jointly trained by the CTC loss and the cross-entropy loss. Experiments on the latest Chinese handwritten text datasets (the SCUT-HCCDoc and the SCUT-EPT) and one English handwritten text dataset (the IAM) show that the proposed method can make a new milestone.
Index Terms:
Handwritten text recognition, segmentation-free recognition, 3D attention module, global-local context information and multi-scale training.
I Introduction
Handwritten text recognition (HTR) is a typical sequence-to-sequence problem. It can be formulated as a Bayesian decision problem. The research efforts include segmentation-based methods [1, 2] and segmentation-free methods [3, 4, 5]. The former methods usually depend on the detected boxes of characters or extra training data. Compared with the segmentation-based methods, only text-level labels are needed during the training stage in the segmentation-free methods.
There are three typical segmentation-free methods, i.e., hidden Markov model (HMM) [3, 6], connectionist temporal classification (CTC)[7, 5, 8, 9, 10] and encoder–decoder (ED) framework [11, 12, 13, 14]. As shown in Fig. 1, in the HMM-based method, each character is modeled by an HMM and a text line can be represented by cascaded HMMs. A series of frames extracted from an original image by a left-to-right sliding window are assigned to the underlying states. Then, a neural network is used to estimate the posterior probabilities of the states, while the outputs of networks in CTC and ED-based approaches are character classes. In the CTC loss, a special character ”blank” and a sophisticated rule are designed to split different characters, and the forward-backward algorithm can efficiently compute the probability of the underlying character sequence. In an encoder-decoder network, an image is usually fed into the encoder to generate the corresponding middle features. Based on the middle representations, the decoder is used to locate and predict the character sequence via attention mechanisms. The common cross-entropy loss can be directly used to adjust the parameters of the ED network.
@@ I Introduction
Although characters can be represented by a high resolution and compact HMM [6], the network used to model the state posterior probability has a large number of output nodes and can not be trained in an end-to-end way. Moreover, it is reasonable to explicitly model the 2D information of characters. However, it is very difficult to expand 1D HMM to 2D HMM [15] due to the computational complexity. Even for the CTC, ED-based approaches [16, 17, 8, 12], early networks just simply depend on the local receptive field of convolutional layers or recurrent units and gradually shrink the height of feature maps to 1 pixel via stacked pooling layers [18], which may lead to information loss.
In this paper, inspired by the above three modeling ways, we propose a new recognition network by using a novel 3D attention module and global-local context information. The 3D attention module is employed to explicitly extract 2D information of blocked feature maps with different resolutions. In detail, the 3D attention module is ingeniously decoupled into a 2D self-attention operation and a 1D attention-based aggregation operation. Based on the outputs (visual features) of the 3D attention module, we further extract the corresponding global-local context information via the self-attention and the recurrent unit, respectively. Finally, by integrating the visual features and the corresponding global-local context features, a well-designed representation for the recognition task can be obtained. In summary, the main contributions of this paper are as follows:
@@
The remainder of this paper is organized as follows: Section II reviews the related work. Section III elaborates on the details of the proposed method. Section IV reports the experimental results and analyses. Finally, we conclude the paper.
II Related Work
In this section, we review related work, including recent advances of the text recognition task, different attention mechanisms and auxiliary features used in the text recognition.
II-A Recent Advances of the Text Recognition Task
The text recognition task includes the scene text recognition and the handwritten text recognition. In recent years, state-of-the-art approaches have changed from multiple procedures to end-to-end training and inference. For the scene text recognition, contrastive learning methods of different levels (character [19, 20], subword and word [21]) have been introduced. Zheng et al. [22] propose the multiplexed routing network for multilingual text recognition. Different from the early attention mechanisms [23, 24], glyph information obtained by the k-means cluster is used to generate more accurate attention. On the other side, many researchers focus on integrating visual and language information [25, 26, 27].
For the handwritten text recognition, Hoang et al. [9] combine the radical-level CTC loss and the character-level loss while Ngo et al. [14] construct a joint decoder to integrate the visual feature and the linguistic context feature. More recently, Peng et al. [2] built a full convolution network to simultaneously achieve the purpose of the character location, the character bounding boxes prediction and the character prediction. Lin et al. [28] propose a mobile text recognizer via searching the lightweight neural units. Furthermore, Peng et al. [29] detect and recognize characters in page-level handwritten text while Coquenet et al. [30] employ a fully convolutional network as the encoder with a transformer decoder for document recognition.
II-B Attention Mechanism
Since the attention was first introduced in the neural machine translation [11], it has been widely used in different fields. The attention methods used in the text recognition task include two aspects: local attention and non-local attention. The former ones can generate a good representation for the current time step, e.g., the channel-wise attention [31], the spatial-wise attention [32], the activation-wise attention [33], the multi-aspect attention [34, 10] in convolutional layers. The non-local attention across different time steps $\bf{X}=\{\bf{x}_{1},...,\bf{x}_{t},...,\bf{x}_{n}\}$can be formulated via three vectors, i.e., the query $\bf{q}_{t}$, the value $\bf{v}_{t}$, and the key $\bf{k}_{t}$ [35]. In a typical self-attention block, the queries, keys and values are transformed from the corresponding input $\bf{X}$, and then The output $\bf{Y}$ is computed as a weighted sum of the values, where the weight assigned to each value is computed by a compatibility function of the query with the corresponding key:
${\bf{Y}}={\rm{softmax(}}\frac{{{\bf{Q}}{{\bf{K}}^{T}}}}{s}{\rm{)}}\bf{V}$ (1)
The matrix ${\bf{Q}}$ packs all query vectors, and all keys and values are also packed together into matrices ${\bf{K}}$ and ${\bf{V}}$. The hyperparameter $s$ is a scaling factor. Most transformer-based recognizers [21, 19, 27] stack the self-attention blocks in the encoder while the previously decoded results are used as the queries in the decoder [25, 26, 34, 20, 30]. In the decoding stage, other effective approaches [24, 13] define the last decoded result as the query vector and assign the corresponding key and value vectors to the same. Specifically, At the time step $t$, the weight coefficient $\alpha_{t,i}$ of the $i$-th value is computed as follows:
${e_{t,i}}$ $={{\bf{w}}^{\rm{T}}}{\rm{tanh}}({\bf{W}}{{\bf{q}}_{t-1}}+{\bf{U}}{{\bf{k}}_{i}}+{\bf{b}})$ (2)
${\alpha_{t,i}}$ $=\frac{{{e_{t,i}}}}{{\sum\limits_{i}{{e_{t,i}}}}}$ (3)
$\bf{w},\bf{W},\bf{U},\bf{b}$ are trainable weights. Similarly, the output $\bf{y}_{t}$ can be obtained:
${\bf{y}_{t}}={\sum\limits_{i}{{\alpha_{t,i}}\bf{v}_{i}}}$ (4)
II-C Auxiliary Features
In the HTR task, an early attempt [36] performs page style clustering by using style features (such as contour slope, pen pressure and writing velocity), and then the authors build independent HMMs for each cluster. With the success of deep learning in a wide range of applications, Wang et al. [6] combine each convolutional layer with one adaptive layer fed by a writer-dependent vector for improving recognition performance. More recently, a well-designed style extractor network trained by identification loss is introduced to explicitly extract personalized writer information [10].
Different from the style features that only exist in the handwritten text, semantic information (context information) is widely used in all text recognition tasks. Many works [3, 1, 37, 38, 6] separately build a character model and a language model or a specialized lexicon. In the decoding stage, by using the language model, the lattices generated by the character model are re-scored to obtain the most reasonable result. Recently, The visual features and the linguistic knowledge were simultaneously integrated into a network [39, 25, 26]. Actually, based on the visual features extracted from a certain image, the corresponding context information can also be naturally obtained by using recurrent units [18, 23, 24, 10] or transformers [40, 30, 27].
III Methodology
@@
As shown in Fig. 2, the proposed network includes three parts, i.e., the convolutional neural network (CNN), the 3D attention module and the features integration block. In the CNN, the hybrid attention module (HAM) [10] is used and each convolutional layer is equipped with the bath normalization [41]. For the outputs of the CNN, we employ a multi-scale framing strategy to extract different solutions. In this section, we elaborate on the details of the 3D attention module and the global-local context information. Moreover, the training pipeline based on the CTC loss and the CE loss is also shown.
@@
Figure 2: The parts within the dosh lines are only used during the training stage. The details of the 3D attention, global-local context, decoder with attention, CE loss and CTC loss are illustrated in the following sections.
III-A 3D Attention
@@
where $i$ is indice along hidden dimensions. And then, we reshape each 3D block from ${C\times S\times H}$ to $SH\times C$ and conduct a self-attention operation for these sequential vectors ${\bf{f}}_{\rm{p}}$ as Eq.1:
@@
Similar to the transformer encoder, the vectors ${\bf{f}}_{p}^{{}^{\prime}}$ are fed into a layer normalization (Layer Norm) followed by two fully connected layers (FCs) before computing Eq.10-12.
@@
III-B Features Integration
Through the 3D attention module, a series of 3D feature maps can be transformed into the corresponding visual features ${\bf{r}}_{t}(t=0...T)$. As shown in Fig. 4-(a), for the visual feature ${\bf{r}}_{t}$, the global context feature ${\bf{l}}_{t}$ and the local context ${\bf{s}}_{t}$ can be extracted by using the self-attention mechanism and the recurrent units, respectively. In this paper, we directly employ the computation method of the ${\bf{f}}^{{}^{\prime}}$ in Fig. 3 to obtain the global features ${\bf{l}}_{t}$. For the local features, the standard LSTM (Fig. 4-(b)) is used. The LSTM consists of forget, input and output gates for maintaining its state over time:
${{\bf{g}}_{t}}$ $={\rm{sigm}}({{\bf{W}}_{f}}{{\bf{r}}_{t}}+{{\bf{U}}_{f}}{{\bf{s}}_{t-1}}+{{\bf{b}}_{f}})$ (13)
${{\bf{i}}_{t}}$ $={\rm{sigm}}({{\bf{W}}_{i}}{{\bf{r}}_{t}}+{{\bf{U}}_{i}}{{\bf{s}}_{t-1}}+{{\bf{b}}_{i}})$ (14)
${{{\bf{\tilde{c}}}}_{t}}$ $={\rm{tanh}}({{\bf{W}}_{\widetilde{c}}}{{\bf{r}}_{t}}+{{\bf{U}}_{\widetilde{c}}}{{\bf{s}}_{t-1}}+{{\bf{b}}_{\widetilde{c}}})$ (15)
${{\bf{c}}_{t}}$ $={{\bf{g}}_{t}}\odot{{\bf{c}}_{t-1}}+{{\bf{i}}_{t}}\odot{{{\bf{\tilde{c}}}}_{t}}$ (16)
${{\bf{o}}_{t}}$ $={\rm{sigm}}({{\bf{W}}_{o}}{{\bf{r}}_{t}}+{{\bf{U}}_{o}}{{\bf{s}}_{t-1}}+{{\bf{b}}_{o}})$ (17)
${{\bf{s}}_{t}}$ $={{\bf{o}}_{t}}\odot\tanh({{\bf{c}}_{t}})$ (18)
where ${\bf{r}}_{t}$ represents the input at time $t$ and ${\bf{s}}_{t}$ is the corresponding output, $\bf{i}$, $\bf{g}$, $\bf{o}$ and $\bf{c}$ are the input gate, forget gate, output gate and cell vectors, respectively. The weight matrix subscripts have the meaning suggested by their names. Finally, the visual feature ${\bf{r}}_{t}$, the global feature ${\bf{l}}_{t}$ and the local feature ${\bf{s}}_{t}$ are simply contacted together:
@@ III-B Features Integration
III-C Joint Training of the CTC and the CE Losses
Given the feature sequence $\bf{V}$ of a text line image, the text recognition task is to find the corresponding underlying n-character sequence $\bf{C}=\{C_{1},C_{2},...,C_{n}\}$, i.e, compute the posterior probability $p(\bf{C}|\bf{V})$. The CTC can be regarded as a loss function of neural networks:
@@ III-C Joint Training of the CTC and the CE Losses
Finally, Alg.1 describes the joint training pipeline.
@@
IV Experiments
IV-A Datasets and Evaluation Metrics
TABLE I: The number of text lines, characters and classes in the SCUT-HCCDoc and the SCUT-EPT.
@@ IV-A Datasets and Evaluation Metrics
TABLE II: The standard partition of the dataset IAM.
@@
Table I lists the number of text lines, characters and classes in both Chinese datasets. In the SCUT-EPT dataset, we removed 681 training text images including abnormal structures, i.e. characters swapping and overlap. Similarly, in the SCUT-HCCDoc dataset, some low-quality images were removed and only 91,261 text instances were used in the training set. All test images in both datasets were evaluated. Table II lists the partition of the English dataset  [^1] https://fki.tic.heia-fr.ch/databases/iam-handwriting-database. All training images, validation images and test images were used in our experiments and the best model was selected for testing according to the results on the validation sets. The character accuracy rate (AR) was used to evaluate the performance of the text recognition model. The evaluation criterion is defined as follows:
@@
IV-B Experimental Results and Analysis
During the training stage, common data argument methods including text location shift, image blur, grey level and contrast changes, and linear transformation were used. The Adam optimization algorithm with default parameters was adopted, we adjusted the learning rate according to training steps until the model converged into a small range. At certain epochs, the learning rate was multiplied by 0.5. The deep learning platform Pytorch [46] and an NVIDIA RTX A6000 with 48GB memory were used.
Firstly, we conducted ablation experiments to verify the effectiveness of the proposed network.
IV-B1 Ablation Experiments
In this section, we examine several important factors for recognition performance. These factors include the frame resolution, the 3D attention module, the global-local context information and the joint training.
We only compare the results of different frame lengths and all sliding windows have the same height. Table III shows the different results of different frame resolutions. We can observe that frame length 3 is the optimal configuration for all experiments, which means the corresponding window size is suitable for most characters. For a longer or shorter feature sequence, it is difficult to discriminate different regions of characters. For example, when we used a small frame length, the Chinese recognition results obviously decreased from 88.8% to 88.06%, 76.67% to 75.13%, respectively, while the English recognition performance only changed from 94.17% to 94.11%, which may be owing to a smaller width for most English characters. Furthermore, we can obtain consistently improvements by integrating multiple resolutions, i.e., the ARs achieve 89.16%, 76.96% and 94.37%, respectively. As shown in Fig. 2, in the training stage, although three 3D attention blocks and global-local context blocks were used in parallel, we only retained the corresponding branch of the frame length 3 during the inference stage.
TABLE III: The different results of different frame resolutions.

### v2

# Improving Handwritten Text Recognition via 3D Attention and Multi-Scale Training

@@ Abstract
The segmentation-free research efforts for addressing handwritten text recognition can be divided into three categories: connectionist temporal classification (CTC), hidden Markov model and encoder–decoder methods. In this paper, inspired by the above three modeling methods, we propose a new recognition network by using a novel three-dimensional (3D) attention module and global-local context information. Based on the feature maps of the last convolutional layer, a series of 3D blocks with different resolutions are split. Then, these 3D blocks are fed into the 3D attention module to generate sequential visual features. Finally, by integrating the visual features and the corresponding global-local context features, a well-designed representation can be obtained. Main canonical neural units including attention mechanisms, fully-connected layer, recurrent unit and convolutional layer are efficiently organized into a network and can be jointly trained by the CTC loss and the cross-entropy loss. Experiments on the latest Chinese handwritten text datasets (the SCUT-HCCDoc and the SCUT-EPT) and one English handwritten text dataset (the IAM) show that the proposed method can achieve comparable results with the state-of-the-art methods. The code is available at https://github.com/Wukong90/3DAttention-MultiScaleTraining-for-HTR.
Keywords:
Handwritten text recognition Segmentation-free recognition 3D attention module Multi-scale training Global-local context information.
1 Introduction
There are three typical segmentation-free methods, i.e., hidden Markov model (HMM) [1, 2], connectionist temporal classification (CTC)[3, 4, 5, 6, 7, 8] and encoder–decoder (ED) framework [9, 10, 11, 12]. As shown in Fig. 1, in the HMM-based method, each character is modeled by an HMM and a text line can be represented by cascaded HMMs. A series of frames extracted from an original image by a left-to-right sliding window are assigned to the underlying states. Then, a neural network is used to estimate the posterior probabilities of the states, while the outputs of networks in CTC and ED-based approaches are character classes. In the CTC loss, a special character "blank" and a sophisticated rule are designed to split different characters, and the forward-backward algorithm can efficiently compute the probability of the underlying character sequence. In an encoder-decoder network, an image is usually fed into the encoder to generate the corresponding middle features. Based on the middle representations, the decoder is used to locate and predict the character sequence via attention mechanisms. The common cross-entropy loss can be directly used to adjust the parameters of the ED network.
@@ 1 Introduction
It is reasonable to explicitly model the 2D information of characters. However, it is very difficult to expand 1D HMM to 2D HMM [13] due to the computational complexity. Even for the CTC, ED-based approaches [14, 15, 6, 10], early networks just simply depend on the local receptive field of convolutional layers or recurrent units and gradually shrink the height of feature maps to small pixels via stacked pooling layers, which may lead to information loss. Although some attempts [16, 17, 11, 18] explicitly or implicitly employ 2D information in scene text recognition and character recognition, there is no obviously evidence shows that these method can be directly used in long handwritten text recognition. Obviously, the context information is important for the text recognition. In [19, 20, 21], the visual features and the linguistic knowledge are simultaneously integrated into a network. Actually, based on the visual features extracted from a certain image, the corresponding context information can also be naturally obtained by using recurrent units [5, 16, 22, 8] or transformers [17, 23, 24].
For the handwritten text recogniton task, recently, Hoang et al. [7] combined the radical-level CTC loss and the character-level loss while Ngo et al. [12] constructed a joint decoder to integrate the visual feature and the linguistic context feature. More recently, Peng et al. [25] built a full convolution network to simultaneously achieve the purpose of the character location, the character bounding boxes prediction and the character prediction. Lin et al. [26] propose a mobile text recognizer via searching the lightweight neural units. Furthermore, Peng et al. [27] detect and recognize characters in page-level handwritten text while Coquenet et al. [23] employ a fully convolutional network as the encoder with a transformer decoder for document recognition.
In this paper, inspired by the above three modeling ways, we propose a new recognition network by using a novel 3D attention module and global-local context information. The 3D attention module is employed to explicitly extract 2D information of block features with different resolutions. In detail, the 3D attention module is ingeniously decoupled into a 2D self-attention operation and a 1D attention-based aggregation operation. Based on the outputs (visual features) of the 3D attention module, we further extract the corresponding global-local context information via the self-attention and the recurrent unit, respectively. Finally, by integrating the visual features and the corresponding global-local context features, a well-designed representation for the recognition task can be obtained. In summary, the main contributions of this paper are as follows:
@@
The remainder of this paper is organized as follows: Section 2 elaborates on the details of the proposed method. Section 3 reports the experimental results and analysis. Finally, we conclude the paper.
2 Methodology
@@
As shown in Fig. 2, the proposed network includes three parts, i.e., the convolutional neural network (CNN), the 3D attention module and the features integration block. In the CNN, the hybrid attention module (HAM) [8] is used and each convolutional layer is equipped with the batch normalization [28]. For the outputs of the CNN, we employ a multi-scale framing strategy to extract different solutions. In this section, we elaborate on the details of the 3D attention module and the global-local context information. Moreover, the multi-scale training pipeline is also shown.
@@
Figure 2: The parts within the dashed lines are only used during the training stage. The details of the 3D attention, global-local context, decoder with attention, CE loss and CTC loss are illustrated in the following sections.
2.1 3D Attention
@@
where $i$ is indice along hidden dimensions. And then, we reshape each 3D block from ${C\times S\times H}$ to $SH\times C$ and conduct a self-attention operation for these sequential vectors ${\bf{f}}_{p}$ as Eq.5:
@@
Similar to the transformer encoder, the vectors ${\bf{f}}_{p}^{{}^{\prime}}$ are fed into a layer normalization (Layer Norm) followed by two fully connected layers (FCs) before computing Eq.6-8.
@@
2.2 Features Integration
Through the 3D attention module, a series of 3D feature maps can be transformed into the corresponding visual features ${\bf{r}}_{t}(t=0...T)$. As shown in Fig. 4-(a), for the visual feature ${\bf{r}}_{t}$, the global context feature ${\bf{l}}_{t}$ and the local context ${\bf{s}}_{t}$ can be extracted by using the self-attention mechanism and the recurrent units, respectively. In this paper, we directly employ the computation method of the ${\bf{f}}^{{}^{\prime}}$ in Fig. 3 to obtain the global context information ${\bf{l}}_{t}$. For the local context features, the standard LSTM (Fig. 4-(b)) is used. ${\bf{r}}_{t}$ represents the input at time $t$ and ${\bf{s}}_{t}$ is the corresponding output, $\bf{i}$, $\bf{g}$, $\bf{o}$ and $\bf{c}$ are the outputs of input gate, forget gate, output gate and cell vectors, respectively. Finally, the visual feature ${\bf{r}}_{t}$, the global feature ${\bf{l}}_{t}$ and the local feature ${\bf{s}}_{t}$ are simply contacted together:
@@ 2.2 Features Integration
2.3 Multi-Scale Training
The proposed multi-scale training include two parts. Firstly, as shown in Fig. 2, in the training stage, multiple feature sequences with different frame lengths are extracted. Although parallel branches including the same 3D attention block and global-local context block are used, we only retained the corresponding branch of the frame length 3 during the inference stage in our experiments.
Secondly, we employ the joint training of the CTC and the CE losses. Given the feature sequence $\bf{V}$ of a text line image, the text recognition task is to find the corresponding underlying n-character sequence $\bf{C}=\{C_{1},C_{2},...,C_{n}\}$, i.e, compute the posterior probability $p(\bf{C}|\bf{V})$. The CTC can be regarded as a loss function of neural networks:
@@ 2.3 Multi-Scale Training
Finally, Alg.1 describes the joint training of the CTC and the CE Losses.
@@
3 Experiments
3.1 Datasets and Evaluation Metrics
Table 1: The number of text lines, characters and classes in the SCUT-HCCDoc and the SCUT-EPT.
@@ 3.1 Datasets and Evaluation Metrics
Table 2: The standard partition of the dataset IAM.
@@
Table 1 lists the number of text lines, characters and classes in both Chinese datasets. In the SCUT-EPT dataset, we removed 681 training text images including abnormal structures, i.e. characters swapping and overlap. Similarly, in the SCUT-HCCDoc dataset, some low-quality images were removed and only 91,261 text instances were used in the training set. All test images in both datasets were evaluated. Table 2 lists the standard partition of the English dataset  [^1] https://fki.tic.heia-fr.ch/databases/iam-handwriting-database. All training images, validation images and test images were used in our experiments and the best model was selected for testing according to the results on the validation sets. The character accuracy rate (AR) was used to evaluate the performance of the text recognition model. The evaluation criterion is defined as follows:
@@
3.2 Experimental Results and Analysis
During the training stage, common data argument methods including text location shift, image blur, grey level and contrast changes, and linear transformation were used. The Adam optimization algorithm with default parameters was adopted, we adjusted the learning rate according to training steps until the model converged into a small range. At certain epochs, the learning rate was multiplied by 0.5. The deep learning platform Pytorch [34] and an NVIDIA RTX A6000 with 48GB memory were used. Firstly, we conducted ablation experiments to verify the effectiveness of the proposed network.
3.2.1 Ablation Experiments
In this section, we examine several important factors for recognition performance. These factors include the frame resolution, the 3D attention module, the global-local context information and the multi-scale training.
We only compare the results of different frame lengths and all sliding windows have the same height. Table 3 shows the different results of different frame resolutions. We can observe that frame length 3 is the optimal configuration for all experiments, which means the corresponding window size is suitable for most characters. For a longer or shorter feature sequence, it is difficult to discriminate different regions of characters. For example, when we used a small frame length, the Chinese recognition results obviously decreased from 88.8% to 88.06%, 76.67% to 75.13%, respectively, while the English recognition performance only changed from 94.17% to 94.11%, which may be owing to a smaller width for most English characters. Furthermore, we can obtain consistently improvements by integrating multiple resolutions, i.e., the ARs achieve 89.16%, 76.96% and 94.37%, respectively. As shown in Fig. 2, in the training stage, although three 3D attention blocks and global-local context blocks were used in parallel, we only retained the corresponding branch of the frame length 3 during the inference stage.
Table 3: The different results of different frame resolutions.